OpenStudy (lilg132):

use integration to evaluate the result of :

OpenStudy (lilg132):

\[{Z = \int\limits_{}^{}7\sin3\theta d \theta} \] at \[ ({\pi \over 6},2)\]

OpenStudy (turingtest):

u-substitution\[u=3\theta\implies du=3d\theta\]

OpenStudy (turingtest):

\[d\theta=\frac73\]so the integral becomes\[\frac73\int\sin udu\]